use power reducing formulas to rewrite the expression in terms of the first power of the cosin.

tan²x
1
Expert's answer
2011-12-20T12:18:15-0500
tan^2(x)=sin^2(x)/cos^2(x)=(1-cos^2(x))/cos^2(x)=1/cos^2(x) -1=2/(1+cos(2x))
-1